OpenStudy (anonymous):

if f(x)=0.5(3-x), what is the value of f(-3)

jimthompson5910 (jim_thompson5910):

replace every copy of x with -3 and evaluate f(x)=0.5(3-x) f(-3)=0.5(3-(-3)) I'll let you finish